OpenStudy (anonymous):

A solid gold cube with edge of 2" is melted down and recast as a cone with a 3" radius base. Find the height of the cone.

OpenStudy (anonymous):

Relplace r in the following equation with 3 and then solve for h.\[1/3\pi r^2h= 2^3\]\[h=\frac{8}{3 \pi } \]

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Thank you for the medal.